OpenStudy (anonymous):

Linda's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Linda $4.10 per pound, and type B coffee costs $5.15 per pound. This month, Linda made 174 pounds of the blend, for a total cost of $795.30. How many pounds of type A coffee did she use?

OpenStudy (anonymous):

I think its called a half and half.

OpenStudy (anonymous):

So I believe the right equation is 4.10(a) + 5.15(b) = 795. a + b = 174

OpenStudy (anonymous):

@ryryfofo

OpenStudy (anonymous):

how do you find a and b?

OpenStudy (anonymous):

So you are gonna want to plug in numbers that add up to 174 into the equation and then find the one that equals 795.30 For example we'll say a = 100 and b = 74 4.10(100) + 5.15(74)= 795 410 + 381.1 = 791.1 410 + 381.1 ≠ 795 however we are close now

OpenStudy (anonymous):

So now try maybe 94 = a, 80 = b

OpenStudy (anonymous):

@ryryfofo

OpenStudy (anonymous):

its not right :/

OpenStudy (anonymous):

Did you get around 766

OpenStudy (anonymous):

Im sorry I used the wrong equation x = pounds type A y = pounds type B Total pounds 1x + 1y = 174 Total value 5.15x + 4.10y = 795.30

OpenStudy (anonymous):

4.10x + 5.15y = 795.30 sorry

OpenStudy (anonymous):

Now you want to multiply the first equation by 5.15 5.15x + 5.15y = 896.1

OpenStudy (anonymous):

Now subtract the first equation from the second one 4.10x + 5.15y = 795.30 -5.15x - 5.15y = -896.1 = -1.05x = -100.8

OpenStudy (anonymous):

Now divide both sides by -1.05 so you get x by itself

OpenStudy (anonymous):

Say something if your still here. Just making sure

OpenStudy (anonymous):

@ryryfofo

OpenStudy (anonymous):

I gave up lol

OpenStudy (anonymous):

x = 96

OpenStudy (anonymous):

oh....

OpenStudy (anonymous):

Sorry... Do you just want the answer then?

OpenStudy (anonymous):

no it's okay

OpenStudy (anonymous):

a = 96, b = 78 the answer is 96 pounds of coffee a

OpenStudy (anonymous):

Medal @ryryfofo